The publisher wants to determine the relationship between the number of copies of books sold and the expenditure made. A sample of 10 books studied shows the following data.

Copies sold
(in thousands) –x
5
13
24
35
58
73
98
100
121
189

Advertising Expense (thousands of pesos)-y
5
9
5
7
8
6
12
21
38
30


1. Determine the regression model.
2. How many copies will be sold if expenditure is P40k ?
3. If there are 220k copies, how much will be the advertising expense?
4. Test the regression coefficient.
1
Expert's answer
2020-12-16T17:58:44-0500

Solution

"\sum x = 716 \space \sum y=141 \space n=10"


"\sum x^2 = 80654 \space \sum y^2 =3209 \sum xy =14953"

1. Regression Model


"\hat y= \hat\alpha +\hat\beta x"

"\hat\beta = {n \sum xy - \sum x \sum y \over n \sum x^2 - {( \sum x)} ^2}"

"={10(14953)-(716*141) \over 10(80654)-{(716)}^2}""=0.1653"

"\alpha = {\sum y - \beta \sum x \over n}"

"={141 - 0.1653(716) \over 10} =2.2645"

2. When expenditure is p40k

y=40


"40=2.2645 +0.1653x"

"x= {40-2.2645 \over 0.1653} = 228.28"

"\approx 228.28k \space copies"


3. If there are 220k copies.

x=220


"y=2.2645 + 0.1653(220) = 38.63"

=p38.63k


4. Test regression coefficient

"H_0 : \beta =0 \space vs \space H_1: \beta \not = 0"


"t={ \beta - 0 \over {S \over \sqrt {SS_x}}}"

"S= \sqrt {SS_y - \beta SS_x \over n-2}"

"SS_y = {n \sum y^2 - {( \sum y)} ^2 \over n}=1220.9"

"SS_x = {n \sum x^2 - {( \sum x)} ^2 \over n} = 29388.4"

"S_{xy} = {n \sum xy - \sum x \sum y \over n} =4857.4"

"S= \sqrt {1220.9 - 0.1653 (4857.4) \over 8} = 7.228"

"t= {0.1653 - 0 \over {7.228 \over 29388.4}} = 3.921"

T-critical

"t_{0.025,8}=2.306"

Since t-stat > t-critical, we reject H0

Therefore the regression coefficient is significant